OpenStudy (anonymous):

graph the function y=lx+2l-3

OpenStudy (usukidoll):

due to the absolute value sign we have two situations. We have to graph y = x+2-3 y=x-1 and y=-(x+2)-3 y=-x-2-3 y=-x-5

OpenStudy (usukidoll):

For any real number x the absolute value or modulus x is denoted by \[\mid x \mid \] (a vertical bar on each side of the quantity) and is defined as \[\mid x \mid = \] x if \[x \geq 0 \] -x if x <0

OpenStudy (anonymous):

I figured it out :)

OpenStudy (usukidoll):

that's good :)